OpenStudy (anonymous):

What is the vertex of the graph of y = – |4x + 8| + 3? (–2, 3) (2, –3) (2, 11) (–2, 5)

Parth (parthkohli):

http://wolframalpha.com I like to cheat it cheat it

OpenStudy (anonymous):

parth no -_-

OpenStudy (anonymous):

I need to learn.........cheating gonna make me lazy

Parth (parthkohli):

Well... the only thing I do here is to use Calculus to find our absolute extrema. Do you want to do that?

OpenStudy (anonymous):

omg big words :O..........sure

Parth (parthkohli):

Do you know about differentiation?

OpenStudy (anonymous):

nope

Parth (parthkohli):

Then you should cheat. And by the way, it is stupid to not use technology when it is there.

Parth (parthkohli):

I'm actually not sure you can use Calculus here. Let me leave it to the experts.

OpenStudy (phi):

equations with absolute values (the | | ) have a vertex where the stuff inside the | | is zero. so find x so that 4x+8 =0 of course the y value is just 3 at this point

OpenStudy (anonymous):

At x=-2 the absolute part is zero; so that is the vertex (-2,3)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thanks for showing your work

OpenStudy (phi):

If you knew the vertex happens when the absolute zero part is 0, they you would see y = -|0| +3 which is just 3 that is y=3 looking at your choices, only 1 pair has 3 for its y-value.

OpenStudy (phi):

If more than 1 pair had 3 for the y value, you would have to find the x value to get the correct pair. that is when you have to solve 4x+8=0 to find 4x= -8 and x= -8/4= -2